OpenStudy (idek-ok):

How do I find the range of a rational function? f(x)=(2)/(x^2-2x-3) PLEASE DON'T ANSWER, JUST AN EXPLANATION OF HOW I WOULD DO IT

OpenStudy (thatonegirl_):

Range is your set of y-values (f(x)) that you get when you plug in x values within the domain.

OpenStudy (thatonegirl_):

\[f(x)=\frac{ 2 }{ x^2-2x-3 }\]

OpenStudy (idek-ok):

Yes, I know. So how would I find that in this situation? I watched a couple of videos, and they said to find the domain of the inverse of the function, but the way they found the inverse does not make sense to me :/

OpenStudy (thatonegirl_):

Well how would you find the domain of this?

OpenStudy (idek-ok):

f(x)=2/x^2-2x3 f(x)=2/(x-3)(x+1) So the domain is (-∞, 3) U (3, ∞), (-∞, -1) U (-1, ∞) right?

OpenStudy (idek-ok):

I struggled with this area as well so I may have just completely missed the mark here. :/

OpenStudy (agent0smith):

You'll likely need to use calculus (derivatives) to find the domain of a non simple rational function.

OpenStudy (idek-ok):

How would I do that? Sorry for the silly question

OpenStudy (agent0smith):

It's not easy to find the domain of the inverse of a rational function like this one... You didn't even answer if you knew calculus...?

OpenStudy (eliesaab):

x=3 is part of the range

OpenStudy (idek-ok):

I am in pre calculus

OpenStudy (agent0smith):

^it's not @eliesaab Well then idk what methods you've learned. You could find the inverse but it requires completing the square or quadratic formula.

OpenStudy (eliesaab):

There are two values for x so that f(x)=3

OpenStudy (eliesaab):

Here they are \[ \left\{\left\{x\to \frac{1}{3} \left(3-\sqrt{42}\right)\right\},\left\{x\to \frac{1}{3} \left(3+\sqrt{42}\right)\right\}\right\} \]

OpenStudy (idek-ok):

The course I am in doesn't really actually explain how to do anything for the most part, which is what is so frustrating thanks for trying to help out!

OpenStudy (eliesaab):

Also there are two values for x such that x=-1 here they are https://www.wolframalpha.com/input/?i=solve+2%2F(x%5E2+-+2+x+-+3)%3D-1

OpenStudy (idek-ok):

Yeah, but I'd like to understand HOW to do it by hand, instead of using a calculator

OpenStudy (eliesaab):

I am tying to tell you that the answer giving above and saying that 3 and -1 are not part of the range is false

OpenStudy (idek-ok):

Oh okay

OpenStudy (agent0smith):

@eliesaab it's clear from the graph the range is not all real numbers

OpenStudy (eliesaab):

In fact the graph tells you that the range is all real numbers except 0

OpenStudy (agent0smith):

You don't see that gap between -0.5 and 0?

OpenStudy (eliesaab):

The answer is \[ (-\infty ,-1/2) \cup (0,\infty) \]

OpenStudy (agent0smith):

I never said anything about 3 or -1, since they're clearly in the range. I was referring to your range of all reals. \[\large y=\frac{ 2 }{ x^2-2x-3 }\] \[\large x=\frac{ 2 }{ y^2-2y-3 }\] \[\large y^2-2y-3 =\frac{ 2 }{ x }\] You can probably use the quad formula or complete the square from here. I am not going to, because this should be done using calculus in the first place.

OpenStudy (eliesaab):

Sorry, the original poster said it.

OpenStudy (agent0smith):

https://www.wolframalpha.com/input/?i=solve+for+y+x%3D+2%2F(y%5E2+-+2y+-+3) You could find the domain from there. But i mean, why? Just either use the graph, or do things like this in calculus.

OpenStudy (eliesaab):

Actually I found a way first take a any number and solve f(x)=a you find two roots, using the quadratic formula Here they are \[ \left\{\left\{x\to \frac{a-\sqrt{2} \sqrt{2 a^2+a}}{a}\right\},\left\{x\to \frac{\sqrt{2} \sqrt{2 a^2+a}+a}{a}\right\}\right\} \] Now \[ 2 a^2 + a\ge 0 \] for a >0 or a <-.5 That is the proof

OpenStudy (eliesaab):

You do not need Calculus here. You only need the quadratic formula

OpenStudy (agent0smith):

It's not that you need calculus, but that it makes it far easier.

OpenStudy (eliesaab):

Actually -.5 is in the range since f(1)=-.5 So finally the answer is \[ (-\infty, -.5]\cup (0,\infty)\]

OpenStudy (eliesaab):

Did I read above that the student is in pre-calc

OpenStudy (agent0smith):

Yeah, but that doesn't change what I said. Finding range of rational functions is much easier with calculus. You won't be able to do much with a more complex rational function (eg. a cubic or higher degree in the denominator)

OpenStudy (sooobored):

wait find the domain? you know that the denominator cannot be equal to zero factor the denominator find the roots and those are the points in the domain that cant exist every other point on the number line is fair game

OpenStudy (sooobored):

oh wait find the range, just look at the extreme points or the value of y as you reach extreme points like as you reach negative infinity, as you reach positive infinity, limit of asymptotes from the left and from the right

OpenStudy (sooobored):

chances are, the y=limit as x-> asymptote might be infinity and the limit as x-> asymptote from the other side might be negative infinity then your range would be (-infinity, infinity) though if all your limits go to 0 and maybe infinity, then you might have to use calculus in order to determine the local minima